When maggie made her monthly trip to the grocery store, she redeemed coupons worth 30 cents, 50 cents, and 55 cents. She used three times as many 30 cent coupons as 55 cent coupons, and four more 50 cent coupons than 55 cent coupons. Maggies's Grocery store doubles the value of all coupons worth up to (but not including) 55 cents. If the total redemption value of Maggie's coupons was $17.40, how many of each coupon did she redeem?
Answer by Paul(988)

Let the 55 cent coupon be x
Let the 30 cent coupon be 3x
Let the 50 cent coupon be x+4
Double = 2(3x)=6x
2(x+4)=2x+8
Total value = 1740
30(6x)+55x+50(2x+8)=1740
180x+55x+100x+400=1740
335x=1340
x=4
4(3)=12
4+4=8

Hence, she has 4, 55 cent coupons, 12, 30 cent coupons and 8, 50 cent coupons.
Paul.